OpenStudy (anonymous):

Find an equation for the line passing through the points (2,4) and (7,-11)

jimthompson5910 (jim_thompson5910):

m=(y2-y1)/(x2-x1) m=(-11-4)/(7-2) m=(-15)/(7-2) m=(-15)/(5) m=-3 So the slope of the line that goes through the points is m=-3 Now let's use the point slope formula: y-y1=m(x-x1) y-4=-3(x-2) y-4=-3x-3(-2) y-4=-3x+6 y=-3x+6+4 y=-3x+10 So the equation of the line that goes through the two points is y=-3x+10

OpenStudy (anonymous):

your sure thats correct eh ??

jimthompson5910 (jim_thompson5910):

what form do they want it in?

OpenStudy (anonymous):

its for math 111 if that makes any sense

jimthompson5910 (jim_thompson5910):

Well the answer is y=-3x+10 if that doesn't work, try placing the answer in standard form which is 3x + y = 10

OpenStudy (anonymous):

thanks alot

OpenStudy (anonymous):

sketch the graphy 3x-4y=24 label x intercpt and y intercept

jimthompson5910 (jim_thompson5910):

plug in x = 0 and solve for y. This gives you one point then plug in y = 0 and solve for x. This gives you the other point to help plot the line

OpenStudy (anonymous):

huh

jimthompson5910 (jim_thompson5910):

so plug in x = 0 3x-4y = 24 3(0)-4y=24 0-4y = 24 -4y=24 y = 24/(-4) y = -6 So the y intercept is (0, -6), which is one point on your graph

OpenStudy (anonymous):

whats the x intercept

jimthompson5910 (jim_thompson5910):

plug in y = 0 3x-4y = 24 3x-4(0) = 24 3x - 0 = 24 3x = 24 x = 24/3 x = 8 So the x intercept is (8,0)